OpenStudy (anonymous):

The slope of the line that goes through (−3, −5) and (−3, −6) is ______ positive negative zero undefined

OpenStudy (cwrw238):

slope = -5-(-6) / -3-(-3)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ -5}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ -6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

ok so yes I know that but like wat is the answer

OpenStudy (cwrw238):

well do the calculation what is -3 - (-3) ?

OpenStudy (anonymous):

6 isnt

OpenStudy (cwrw238):

no -(-3) = + 3 so we have -3 + 3 = ?

OpenStudy (anonymous):

-6

OpenStudy (cwrw238):

no -3 + 3 = 3 - 3 = 0

OpenStudy (anonymous):

oo ok sorry

OpenStudy (cwrw238):

so the slope is -6-(-5) / 0 = -1 / 0 - we have a fraction which is a number divided by 0 does it make sense to divide by 0 ??

OpenStudy (anonymous):

ok yes I got now thanx

OpenStudy (cwrw238):

it does not make sense to divide by ero the answer is the slope is undefined

OpenStudy (cwrw238):

the line is vertical